# MoGe: Unlocking Accurate Monocular Geometry Estimation for Open-Domain Images with Optimal Training Supervision

### v1

@@ 4.1 Quantitative Evaluations
We assess the zero-shot performance of our model and compare it with several state-of-the-art methods for monocular point map estimation [69, 41, 57] and depth map estimation [65, 66, 6, 26, 5, 20, 29]. To ensure a fair comparison, all methods utilize ViT-Large [15] as backbone, except for LeReS (which employs ResNeXt101 [63]) as well as GeoWizard and Marigold (both use StableDiffusion V1 [45]). For a valid evaluation, the scale and shift of the predicted point maps and depth maps are first aligned with the ground truth when necessary. We also evaluate the FoV prediction results against several recent methods [28, 75]. As detailed below, our method achieves superior results across all evaluation metrics for various tasks on different benchmarks.
@@
We evaluate the accuracy of point map estimation on eight diverse datasets: NYUv2 [38], KITTI [51], ETH3D [49], iBims-1 [31], Sintel [9], Google Scanned Objects (GSO) [16], DDAD [24], and DIODE [52]. These datasets represent a wide range of domains, including indoor scenes, street views, object scans, and synthetic movies. The raw datasets are processed for reliable evaluation (e.g., sky regions in Sintel dataset and boundary artifacts in the DIODE are removed; see the suppl. material for more preprocessing details). For evaluation metrics, we use the relative point error Relp, i.e., $\|\mathbf{p}-\hat{\mathbf{p}}\|_{2}/\|\hat{\mathbf{p}}\|_{2}$, and the percentage of inliers $\delta_{1}^{\text{p}}$ with $\|\mathbf{p}-\hat{\mathbf{p}}\|_{2}/\|\hat{\mathbf{p}}\|_{2}<0.25$.
We evaluate our method and compare it with LeReS [69], UniDepth [41], and DUSt3R [57]. LeReS and UniDepth predict depth and camera parameters from single images. DUSt3R is designed for two-view input and directly outputs camera-space point maps. It accepts two identical images for handling monocular scenarios. We evaluate both scale-invariant point maps in camera space and affine-invariant point maps. Table 2 demonstrates that our method outperforms existing methods across most benchmarks, achieving the lowest average Relp for both point maps representations.
To further evaluate local geometry accuracy, we assess affine-invariant point maps within local regions defined by object segmentation masks, either from the dataset labels or obtained using Segment Anything [30]. Since per-region point map evaluation demands high-quality ground truth labels as well as the composition of multiple objects in a scene, NYUv2, KITTI, and GSO are excluded. Compared to the previous methods, our method significantly reduces the region-wise Relp by a large margin: from $7.96$ to $5.33$, achieving approximately a $30\%$ error reduction.
@@
To conduct a comprehensive comparison with different monocular depth estimation methods, we evaluate scale-invariant depth [69, 5, 57, 26, 41], affine-invariant depth [29, 20], and affine-invariant disparity [6, 65, 66]. The $z$ coordinates in our predicted point maps represent affine-invariant depth. By utilizing the camera shifts derived from Eq. 2, we convert these coordinates into scale-invariant depth. For affine-invariant disparity, we take the inverse of our scale-invariant depth predictions.
As shown in Table 3, our method achieves the best performance across most benchmarks and exhibits the lowest average Reld for all depth representations. Since existing methods with scale-invariant depth representations can be converted to affine-invariant depth and disparity, we provide an extensive comparison in the suppl. materials, where our method also significantly outperforms the others.
@@
To evaluate the accuracy of the estimated camera FoV, we select benchmarks with vertical $\text{vertical FoV}\geq 45^{\circ}$, including NYUv2 [38], ETH3D [49] and iBims-1 [31], and augment the FoV by randomly center cropping 50$\sim$100from the original images. All input images are undistorted and have a centered principal point.
We compare our model to point map estimation methods [69, 41, 57] as well as two learning-based camera calibration methods: Perspective [28] and WildCam [75]. For fair assessments of zero-shot performance, we utilize their pre-trained checkpoints on in-the-wild datasets. Mean and median FoV errors are reported in Table 5. Our method achieves a mean error of $2.91^{\circ}$ and a median error of $2.21^{\circ}$ on average of the three FoV benchmarks, outperforming both camera calibration methods and point map estimation methods.
@@ 4.2 Qualitative comparison
We show visual results of disparity and point maps of different methods in Fig. 7. For depth prediction methods that do not recover camera parameters, we lift their results to 3D points using ground-truth intrinsics when available. Upon visual inspection, our method produces significantly less-distorted point maps compared to others, demonstrating its superior generalizabiltiy and accuracy. For example, as shown in the fourth row of Fig. 7, other methods result in noticeable stretching and deformation of the car shape. Although [66] produces visually sharp disparity maps, distorted 3D geometry can be observed even when provided with ground-truth camera intrinsics. In contrast, our method yields a more regular and well-formed car structure.
Figure 7: Qualitative comparison of point and disparity maps. *: for methods without camera intrinsics prediction, we use ground-truth camera intrinsics or disparity shifts to lift their results into 3D points when available. The first four cases are samples from our test datasets, while the last two are random “in-the-wild” images.
@@ 4.3 Ablation Study
We conducted comprehensive ablation studies on our point map representation, alignment strategy and loss function designs. For efficiency, we use the ViT-base model as backbone for these experiments.
@@
Our method predicts affine-invariant point maps. Here we compare this representation with two variants, one scale-invariant point map in camera space, and another that predicts a depth maps and camera parameters. For the former, we apply the proposed ROE alignment solver to solve the scale parameter only. For the latter, we adopt the SI-log loss [18] for the depth map and the MSE loss for the ray direction map, similar to UniDepth [41]. The first three rows of Table 5 show the results of different approaches. In these experiments, only the global ground-truth supervision is applied during training. Our proposed affine-invariant representation outperforms the other variants across various tasks. This can be attributed to the fact that the other two variants suffer from focal-distance ambiguity issue during training which jeopardizes their performance.
Optimal alignment.
To evaluate the effectiveness of our optimal alignment, we compare it to a previously-used simple alignment method [42], i.e., we normalize the predicted and ground-truth point maps by their median $z$: $(\mathbf{p}-\mathbf{t}(\mathbf{P}))/s(\mathbf{P})$, where $\mathbf{t}(\mathbf{P})$ equals to $(0,0,\operatorname{median}(\mathbf{Z}))$, and $s(\mathbf{P})$ is computed as $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{p}_{i}-\mathbf{t}(\mathbf{P})\|_{1}$. As shown in the third and fourth rows of Table 5, our proposed alignment method consistently outperforms the median alignment strategy across all tasks by large margins, demonstrating the critical role of our proposed optimal alignment.
@@
Our method predicts a valid region mask to handle regions at infinity. As shown in Fig. 6 (b), our method correctly predicts the sky regions, for which the predicted point values will be erroneous if the masks are removed. We further compare our mask prediction design with another strategy for addressing infinity. Specifically, we assign a large value, i.e., $1,000$ times the average distance of valid regions, to the infinity regions. Fig. 6 (b) shows that this strategy compromises the accuracy of foreground predictions as the network struggles to generate large values for infinity. Incorporating a separate mask prediction offers a convenient solution for masking out infinity and retaining accurate foreground geometry learning.
@@ 5 Conclusion
We have presented a method for accurate monocular geometry estimation of open-domain images. Our key insight is to design effective supervision which is largely neglected by previous methods. We propose a new affine-invariant point map representation which precludes ambiguous supervision during training. A robust, optimal, and efficient point map alignment solver is introduced for accurate global shape learning, together with a multi-scale local geometry loss designed for precise local geometry supervision. Trained on a large dataset collection, our model demonstrates strong generalizability and high accuracy for open-domain images. It significantly outperforms previous methods across various tasks and benchmarks. We believe our method marks a significant advancement in monocular 3D geometry estimation and can serve as a robust foundation model for various vision applications.
@@ A.1 Recovering shift and camera focal
$\min_{f,t_{z}}\sum_{i\in\mathcal{M}}\left({fx_{i}\over z_{i}+t_{z}^{\prime}}-u_{i}\right)^{2}+\left({fy_{i}\over z_{i}+t_{z}^{\prime}}-v_{i}\right)^{2}$ (9)
which can be further reduced to have a single variable $t_{z}^{\prime}$ by substituting $f$ with its close-form solution with respect to $t_{z}^{\prime}$
$f={\sum_{i\in\mathcal{M}}\left(x_{i}\over z_{i}+t_{z}\right)u_{i}+\sum_{i\in\mathcal{M}}\left(y_{i}\over z_{i}+t_{z}\right)v_{i}\over\sum_{i\in\mathcal{M}}\left(x_{i}\over z_{i}+t_{z}\right)^{2}+\sum_{i\in\mathcal{M}}\left(y_{i}\over z_{i}+t_{z}\right)^{2}}$ (10)
@@ A.2 Optimal alignment
Subproblem (w/o clipping)
Consider the optimization objective with respect to scale $s$ only, denoted as $l_{0}(s)$. We omit the mask $\mathcal{M}$ for simplicity and denote $N$ as the number of valid points.
$\min_{s}l_{0}(s)=\min_{s}\sum_{i=1}^{N}w_{i}|s\hat{x}_{i}-x_{i}|$ (11)
@@
$\begin{split}l_{0}^{\prime-}(s)&=\sum_{{x_{i}\over\hat{x}_{i}}<s}{w_{i}\hat{x}_{i}}-\!\!\sum_{s\leq{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}<s}w_{i}\hat{x}_{i}-\sum_{i=1}^{N}w_{i}\hat{x}_{i}\\ l_{0}^{\prime+}(s)&=\sum_{{x_{i}\over\hat{x}_{i}}\leq s}{w_{i}\hat{x}_{i}}-\!\!\sum_{s<{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}\leq s}w_{i}\hat{x}_{i}-\sum_{i=1}^{N}w_{i}\hat{x}_{i}\end{split}$ (12)
$l_{0}^{\prime-}(s)$ and $l_{0}^{\prime+}(s)$ differ at $\left\{x_{i}\over\hat{x}_{i}\right\}$. One might first sort $\left\{x_{i}\over\hat{x}_{i}\right\}$ and compute the prefix summations of $\{w_{i}\hat{x}_{i}\}$. Then the derivatives can be evaluated in $O(1)$ time for each point in $\left\{x_{i}\over\hat{x}_{i}\right\}$. Finally, $\hat{x}_{i}\over x_{i}$ such hat ł$\l_{0}^{\prime-}\left(x_{i}\over\hat{x}_{i}\right)<0\leq\l_{0}^{\prime+}\left(x_{i}\over\hat{x}_{i}\right)$ is a deterministic minimum point.
Subproblem (w/ clipping)
We clip each residual term to suppress the contribution of outliers. Specifically, the clipped objective is
$\min_{s}l_{1}(s)=\min_{s}\sum_{i=1}^{N}\min(\tau,w_{i}|s\hat{x}_{i}-x_{i}|)$ (13)
For each item $l_{1,i}(s)=\min(\tau,w_{i}|s\hat{x}_{i}-x_{i}|)$ in the equation, the one-sided derivatives are
$\begin{split}l_{1,i}^{\prime-}(s)=\left\{\begin{array}[]{cc}-w_{i}\hat{x}_{i}&{w_{i}x_{i}-\tau_{i}\over w_{i}\hat{x}_{i}}<s\leq{x_{i}\over\hat{x}_{i}}\\ w_{i}\hat{x}_{i}&{x_{i}\over\hat{x}_{i}}<s\leq{w_{i}x_{i}+\tau_{i}\over w_{i}\hat{x}_{i}}\\ 0&\text{otherwise}\end{array}\right.\\ l_{1,i}^{\prime+}(s)=\left\{\begin{array}[]{cc}-w_{i}\hat{x}_{i}&{w_{i}x_{i}-\tau_{i}\over w_{i}\hat{x}_{i}}\leq s<{x_{i}\over\hat{x}_{i}}\\ w_{i}\hat{x}_{i}&{x_{i}\over\hat{x}_{i}}\leq s<{w_{i}x_{i}+\tau_{i}\over w_{i}\hat{x}_{i}}\\ 0&\text{otherwise}\end{array}\right.\end{split}$ (14)
@@
$\begin{split}l_{1}^{\prime-}(s)&=\sum_{i=1}^{N}l_{1,i}^{\prime-}(s)\\ &=\!\!\!\!\sum_{{x_{i}\over\hat{x}_{i}}<s\leq{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}}{w_{i}\hat{x}_{i}}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}<s\leq{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}<s}w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}<s}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}<s}\!\!\!\!w_{i}\hat{x}_{i}\end{split}$ (15)
$\begin{split}l_{1}^{\prime+}(s)&=\sum_{i=1}^{N}l_{1,i}^{\prime+}(s)\\ &=\!\!\!\!\sum_{{x_{i}\over\hat{x}_{i}}\leq s<{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}}{w_{i}\hat{x}_{i}}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\leq s<{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}\leq s}w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\leq s}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}\leq s}\!\!\!\!w_{i}\hat{x}_{i}\end{split}$ (16)
One can evaluate $l_{1}^{\prime}(s)$ efficiently through first binary searching on the sorted arrays $\left\{x_{i}\over\hat{x}_{i}\right\}$, $\left\{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}\right\}$ and $\left\{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}\right\}$ and then indexing the precomputed prefix summations of $\{w_{i}\hat{x}_{i}\}$ in the associated orders.
@@
There exists at least one pair of $(k^{*},s^{*})$ such that $s^{*}x_{k^{*}}-\hat{x}_{k^{*}}=0$ and $s^{*}$ minimizes Eq. 13

### v2

@@ 4.1 Quantitative Evaluations
We assess the zero-shot performance of our model and compare it to several state-of-the-art methods for monocular point map estimation [71, 42, 59] and depth map estimation [67, 68, 6, 26, 5, 20, 29]. To ensure a fair comparison, all methods utilize ViT-Large [15] as backbone, except for LeReS (using ResNeXt101 [65]) as well as GeoWizard and Marigold (using StableDiffusion V1 [46]). The scale and shift of the predicted point maps and depth maps are aligned with the ground truth when necessary. We also evaluate the FOV prediction results against several recent methods [28, 77]. As detailed below, our method achieves superior results across various tasks and benchmarks.
@@
We evaluate the accuracy of point map estimation on eight diverse datasets: NYUv2 [39], KITTI [53], ETH3D [50], iBims-1 [31], Sintel [9], Google Scanned Objects (GSO) [16], DDAD [24], and DIODE [54]. These datasets represent a wide range of domains, including indoor scenes, street views, object scans, and synthetic movies. The raw datasets are processed for reliable evaluation (e.g., sky regions in Sintel dataset and boundary artifacts in the DIODE are removed; see the suppl. material for more preprocessing details). For evaluation metrics, we use the relative point error Relp, i.e., $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\|\mathbf{p}\|_{2}$, and the percentage of inliers $\delta_{1}^{\text{p}}$ with $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\|\mathbf{p}\|_{2}<0.25$.
We compare our method with LeReS [71], UniDepth [42], and DUSt3R [59], evaluating both camera-space scale-invariant point maps and affine-invariant point maps. Table 1 demonstrates that our method outperforms existing methods, achieving the lowest average Relp and highest $\delta_{1}^{\text{p}}$ for both point map representations.
To further assess local geometry accuracy, we evaluate affine-invariant point maps within regions defined by object segmentation masks, sourced from dataset labels or Segment Anything [30]. Due to the need for high-quality ground truth and scenes with multiple objects, NYUv2, KITTI, and GSO are excluded. Our method achieves a substantial reduction in region-wise Relp, lowering the error from $7.96$ to $5.33$—an approximately $30\%$ improvement over previous methods.
@@
For a comprehensive comparison with different monocular depth estimation methods, we evaluate scale-invariant depth [71, 5, 59, 26, 42], affine-invariant depth [29, 20], and affine-invariant disparity [6, 67, 68]. The $z$ coordinates in our predicted point maps represent affine-invariant depth, which can be converted to scale-invariant depth using the shift derived from Eq. 2. For affine-invariant disparity, we take the inverse of our scale-invariant depth predictions. As shown in Table 2, our method exhibits the lowest average Reld for all depth representations. Since existing methods with scale-invariant depth representations can be converted to affine-invariant depth and disparity, we provide an extensive comparison in the suppl. materials, where our method also significantly outperforms the others.
@@
To evaluate the accuracy of the estimated camera FOV, we select benchmarks with vertical $\text{vertical FOV}\geq 45^{\circ}$, including NYUv2 [39], ETH3D [50] and iBims-1 [31], and augment the FOV by randomly center cropping 50$\sim$100from the original images. All input images are undistorted with a centered principal point.
We compare our model to point map estimation methods [71, 42, 59] as well as two learning-based camera calibration methods: Perspective Fields [28] and Wild Camera [77]. For fair zero-shot evaluation, we utilize their pre-trained checkpoints on in-the-wild datasets. Table 4 reports mean and median FOV errors. Our method achieves a mean error of $2.91^{\circ}$ and a median error of $2.21^{\circ}$ on average of the three FOV benchmarks, outperforming both camera calibration and point map estimation methods.
@@ 4.2 Qualitative comparison
We show visual results of disparity and point maps of different methods in Fig. 7. For depth prediction methods without camera parameters estimation, we lift their results to 3D points using ground-truth intrinsics when available. Upon visual inspection, our method produces significantly less-distorted point maps compared to others, demonstrating its superior generalizability and accuracy. For example, as shown in the third row of Fig. 7, other methods show noticeable stretching and deformation of the car shape. Although [68] produces visually sharp disparity maps, distorted 3D geometry is evident even with ground-truth camera intrinsics. In contrast, our method yields a more regular and well-formed car structure. More comparisons can be found in the suppl. materials.
Figure 7: Qualitative comparison of point and disparity maps. The first three cases are from our test datasets, while the last two rows are random “in-the-wild” images. *: for methods without camera intrinsics prediction, we use ground-truth camera intrinsics (and disparity shifts) to lift their results into 3D points when available, or use our intrinsics prediction for the “in-the-wild” images. Best view with zoom.
@@ 4.3 Ablation Study
We conducted comprehensive ablation studies on our point map representation, alignment strategy and loss function designs. ViT-base backbones are used here for efficiency.
@@
We compare our affine-invariant point map representation with two variants, one scale-invariant point map in camera space, and another that predicts a depth map and camera parameters. For the former, we apply the proposed ROE alignment to solve the scale parameter only. For the latter, we adopt the SI-log loss [18] for the depth map and the MSE loss for the ray direction map, similar to UniDepth [42]. Results are shown in the first three rows of Table 4, where only the global ground-truth supervision is applied in these experiments. Our proposed affine-invariant representation outperforms the others across various tasks. This is attributed to the fact that the other two variants suffer from focal-distance ambiguity issue during training, which hinders their performance.
ROE alignment.
To evaluate the effectiveness of our optimal alignment, we compare it to a previously-used simple alignment method [43], i.e., the predicted and ground-truth point maps are normalized by their median: $(\mathbf{p}-\mathbf{t}(\mathbf{P}))/s(\mathbf{P})$, where $\mathbf{t}(\mathbf{P})$ equals to $(0,0,\operatorname{median}(\mathbf{Z}))$, and $s(\mathbf{P})$ is computed as $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{p}_{i}-\mathbf{t}(\mathbf{P})\|_{1}$. As shown in the third and fourth rows of Table 4, our proposed alignment method consistently outperforms the median alignment strategy across all tasks by large margins, demonstrating the critical role of our proposed optimal alignment.
We ablate the truncation of residuals in alignment objects, which degrades performance (see the fifth row in Table 4). In practice, adding truncation also improves training stability.
@@
Our method predicts a valid region mask to handle infinity regions. As shown in Fig. 6 (b), it correctly identifies sky regions, where the output point values would be erroneous if the masks were removed. Another strategy for addressing infinity is to assign a large value, i.e., 1,000×1,000$\times$1 , 000 × the average distance, to these regions. Fig. 6 (b) shows that this strategy compromises the foreground accuracy as the network struggles with large values for infinity. Incorporating a separate mask prediction effectively handles infinity while retaining accurate foreground geometry.
@@ 5 Conclusion
We have presented a method for accurate monocular geometry estimation of open-domain images. Our key insight is to design effective supervision which is largely neglected by previous methods. We introduce an affine-invariant point map representation which precludes ambiguous supervision during training. A robust, optimal, and efficient point map alignment solver is introduced for effective shape learning, along with a multi-scale geometry loss for precise local supervision. Trained on a large dataset collection, our model demonstrates strong generalizability and significantly outperforms previous methods across various tasks and benchmarks. We believe our method marks a significant advancement in monocular 3D geometry estimation and can serve as a robust foundational model for various applications.
@@ A.1 Recovering Shift and Camera Focal
$\min_{f,t_{z}}\sum_{i\in\mathcal{M}}\left({fx_{i}\over z_{i}+t_{z}^{\prime}}-u_{i}\right)^{2}+\left({fy_{i}\over z_{i}+t_{z}^{\prime}}-v_{i}\right)^{2},$ (9)
which can be further reduced to have a single variable $t_{z}^{\prime}$ by substituting $f$ with its close-form solution with respect to $t_{z}^{\prime}$,
$f={\sum_{i\in\mathcal{M}}\left(x_{i}\over z_{i}+t_{z}^{\prime}\right)u_{i}+\sum_{i\in\mathcal{M}}\left(y_{i}\over z_{i}+t_{z}^{\prime}\right)v_{i}\over\sum_{i\in\mathcal{M}}\left(x_{i}\over z_{i}+t_{z}^{\prime}\right)^{2}+\sum_{i\in\mathcal{M}}\left(y_{i}\over z_{i}+t_{z}^{\prime}\right)^{2}}.$ (10)
@@ A.2 ROE Alignment
Subproblem (w/o truncation).
Consider the optimization objective with respect to scale $s$ only, denoted as $l_{0}(s)$. We omit the mask $\mathcal{M}$ for simplicity and denote $N$ as the number of valid points:
$\min_{s}l_{0}(s)=\min_{s}\sum_{i=1}^{N}w_{i}|s\hat{x}_{i}-x_{i}|,$ (11)
@@
$\begin{split}l_{0}^{\prime-}(s)&=\!\!\sum_{{x_{i}\over\hat{x}_{i}}<s}{w_{i}\hat{x}_{i}}-\!\!\sum_{s\leq{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}=2\sum_{{x_{i}\over\hat{x}_{i}}<s}w_{i}\hat{x}_{i}-\!\!\sum_{i=1}^{N}w_{i}\hat{x}_{i},\\ l_{0}^{\prime+}(s)&=\!\!\sum_{{x_{i}\over\hat{x}_{i}}\leq s}{w_{i}\hat{x}_{i}}-\!\!\sum_{s<{x_{i}\over\hat{x}_{i}}}{w_{i}\hat{x}_{i}}=2\sum_{{x_{i}\over\hat{x}_{i}}\leq s}w_{i}\hat{x}_{i}-\!\!\sum_{i=1}^{N}w_{i}\hat{x}_{i}.\end{split}$ (12)
$l_{0}^{\prime-}(s)$ and $l_{0}^{\prime+}(s)$ differ at $\{{x_{i}\over\hat{x}_{i}}\}$. First, we sort $\{{x_{i}\over\hat{x}_{i}}\}_{i=1}^{N}$ and compute the prefix summations of $\{w_{i}\hat{x}_{i}\}_{i=1}^{N}$. This allows us to evaluate the derivatives in $O(1)$ time for each point in $\{{x_{i}\over\hat{x}_{i}}\}_{i=1}^{N}$. Finally, $\hat{x}_{i}\over x_{i}$ such that ł$\l_{0}^{\prime-}({x_{i}\over\hat{x}_{i}})\leq 0\leq\l_{0}^{\prime+}({x_{i}\over\hat{x}_{i}})$ is the minimum point. The solution is outlined in Algorithm 2.
Algorithm 2 ROE alignment subproblem w/o truncation
input: arrays $\hat{X}[1..n]$, $X[1..n]$, $W[1..n]$
output: optimal scale $s^{*}$ and objective value $l^{*}$ to Eq. 11
function SolveSubproblem($\hat{X}$, $X$, $W$)
       sort arrays $\hat{X}$, $X$, $W$ by $X[i]/\hat{X}[i]$
       $Q[1..n]$leftarrow accumulated sum of $W*\hat{X}$
       $D[0..n]$leftarrow $\{-Q[n]\}\cup\{2\cdot Q[i]-Q[n]\}_{i=1}^{n}$
       $i^{*}$leftarrow the first $i$ s.t. $D[i-1]\leq 0\leq D[i]$
       $s^{*}$leftarrow $X[i^{*}]/\hat{X}[i^{*}]$
       $l^{*}$leftarrow objective function value at $s^{*}$.
       return $s^{*}$, $l^{*}$.
end function
Subproblem (w/ truncation).
We truncate each residual term to suppress outliers. The truncated objective is
$\min_{s}l_{1}(s)=\min_{s}\sum_{i=1}^{N}\min(\tau,w_{i}|s\hat{x}_{i}-x_{i}|),$ (13)
where $\tau$ is set to $1$ in all our experiments. For each item $l_{1,i}(s)=\min(\tau,w_{i}|s\hat{x}_{i}-x_{i}|)$ in the equation, the one-sided derivatives are
$\begin{split}l_{1,i}^{\prime-}(s)=\left\{\begin{array}[]{cc}-w_{i}\hat{x}_{i}&{w_{i}x_{i}-\tau_{i}\over w_{i}\hat{x}_{i}}<s\leq{x_{i}\over\hat{x}_{i}}\\ w_{i}\hat{x}_{i}&{x_{i}\over\hat{x}_{i}}<s\leq{w_{i}x_{i}+\tau_{i}\over w_{i}\hat{x}_{i}}\\ 0&\text{otherwise}\end{array}\right.,\\ l_{1,i}^{\prime+}(s)=\left\{\begin{array}[]{cc}-w_{i}\hat{x}_{i}&{w_{i}x_{i}-\tau_{i}\over w_{i}\hat{x}_{i}}\leq s<{x_{i}\over\hat{x}_{i}}\\ w_{i}\hat{x}_{i}&{x_{i}\over\hat{x}_{i}}\leq s<{w_{i}x_{i}+\tau_{i}\over w_{i}\hat{x}_{i}}\\ 0&\text{otherwise}\end{array}\right..\end{split}$ (14)
@@
$\begin{split}l_{1}^{\prime-}(s)&=\sum_{i=1}^{N}l_{1,i}^{\prime-}(s)=\!\!\!\!\sum_{{x_{i}\over\hat{x}_{i}}<s\leq{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}}\!\!\!\!{w_{i}\hat{x}_{i}}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}<s\leq{x_{i}\over\hat{x}_{i}}}\!\!\!\!{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}<s}w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}<s}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}<s}\!\!\!\!w_{i}\hat{x}_{i},\end{split}$ (15)
$\begin{split}l_{1}^{\prime+}(s)&=\sum_{i=1}^{N}l_{1,i}^{\prime+}(s)=\!\!\!\!\sum_{{x_{i}\over\hat{x}_{i}}\leq s<{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}}\!\!\!\!{w_{i}\hat{x}_{i}}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\leq s<{x_{i}\over\hat{x}_{i}}}\!\!\!\!{w_{i}\hat{x}_{i}}\\ &=2\sum_{{x_{i}\over\hat{x}_{i}}\leq s}w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\leq s}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}\leq s}\!\!\!\!w_{i}\hat{x}_{i}.\end{split}$ (16)
Lemma 1 shows that the minimum of the objective function in Eq. 13 is still achieved at one of the points in the set $\{{x_{i}\over\hat{x}_{i}}\}_{i=1}^{N}$, despite the function is non-convex and may contain local minima.
Solving the subproblem requires two steps, as outlined in Algorithm 3. The first step is to identify all extrema in $\{{x_{i}\over\hat{x}_{i}}\}_{i=1}^{N}$ that satisfy $l_{0}^{\prime-}({x_{i}\over\hat{x}_{i}})<0\leq l_{0}^{\prime+}({x_{i}\over\hat{x}_{i}})$ by evaluating the derivative values. This can be done efficiently through first binary searching on the sorted arrays $\{{x_{i}\over\hat{x}_{i}}\}_{i=1}^{N}$, $\{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}\}_{i=1}^{N}$ and $\{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}\}_{i=1}^{N}$ and then indexing the prefix summations of $\{w_{i}\hat{x}_{i}\}$ in the associated orders. This step has a complexity of $O(N\log N)$. The second step involves computing the objective values at these extrema and determining the minimum, which takes $O(Nn_{e})$ time, where $n_{e}$ is the number of extrema. As $n_{e}$ approximates the number of outliers, it is typically a small constant in practice.
@@
There exists at least one pair of $(k^{*},s^{*})$ such that $s^{*}\hat{x}_{k^{*}}-x_{k^{*}}=0$ and $s^{*}$ minimizes Eq. 13.